3) Give the formation reaction for rubidium hydrogen carbonate (RbHCO3(S)).
Answer #1

Standard enthalpy of formation is the change of enthalpy during the formation of 1 mole of substance from its constituent elements, with all substances in their standard states.

RuHCO3 has four elements.

Elemental state of Rubidium = Rb (s)

Elemental state of Hydrogen = H2 (g)

Elemental state of Carbon = C (s)

Elemental state of Oxygen = O2 (g)

In the formation reaction, 1 mole of RuHCO3 is formed from its constituent elements.

So, the formation reaction is

Rb (s) + ½ H2(g) + C (s) + 3/2 O2(g) ---------> RbHCO3(s)